### v1

# Lie algebra homology with coefficients tensor products of the adjoint representation in relative polynomial degree $2$

Abstract.
The homology of free Lie algebras with coefficients in tensor products of the adjoint representation working over $\mathbb{Q}$ contains important information on the homological properties of polynomial outer functors on free groups. The latter category was introduced in joint work with Vespa, motivated by the study of higher Hochschild homology of wedges of circles.
There is a splitting of this homology by polynomial degree (for polynomiality with respect to the generators of the free Lie algebra) and one can consider the polynomial degree relative to the number of tensor factors in the coefficients. It suffices to consider the Lie algebra homology in homological degree one; this vanishes in relative degree 00 and is readily calculated in relative degree $1$.
This paper calculates the homology in relative degree $2$, which presents interesting features. This confirms a conjecture of Gadish and Hainaut.
Key words and phrases:
Lie algebra homology; adjoint representation; polynomial functor
2020 Mathematics Subject Classification:
17B01,17B56
1. Introduction
For any Lie algebra $\mathfrak{g}$ (working here over $\mathbb{Q}$) and natural number $r$, one can form the Lie algebra homology $H_{*}(\mathfrak{g};\mathfrak{g}^{\otimes r})$ of $\mathfrak{g}$ with coefficients in the $r$th tensor product of the adjoint representation. The case $r=0$ corresponds to trivial coefficients $\mathbb{Q}$; the case $r=1$ is homology with coefficients in the adjoint representation. Since the action of the symmetric group $\mathfrak{S}_{r}$ by place permutations on $\mathfrak{g}^{\otimes r}$ is by $\mathfrak{g}$-module isomorphisms, $H_{*}(\mathfrak{g};\mathfrak{g}^{\otimes r})$ inherits a $\mathfrak{S}_{r}$-module structure.
Here we are interested in the case where $\mathfrak{g}$ is the free Lie algebra $\mathbb{L}\mathrm{ie}(V)$ on a finite-dimensional $\mathbb{Q}$-vector space $V$, considered as a functor of $V$. This is motivated by the relationship with higher Hochschild homology arising from the author’s joint work with Christine Vespa [PV25]. More explicitly, in that paper we established the relationship with the category of outer functors, namely functors from $\mathbf{gr}$ (the category of finitely-generated free groups) to $\mathbb{Q}$-vector spaces on which inner automorphisms of free groups act trivially. In [Pow24b], using the infinitesimal interpretation of polynomial functors on free groups given in [Pow24a], this was related to Lie algebra homology with coefficients in tensor products of the adjoint representation. This is also related to work of Nir Gadish and Louis Hainaut [GH24], who study the homology of certain configuration spaces and relate this to higher Hochschild homology; their calculations are used by Hainaut in [Hai23] to study homological properties of the category of polynomial outer functors.
Working with the free Lie algebra, the question reduces to that of calculating the representations of $\mathfrak{S}_{r}$
$H_{0}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$
naturally with respect to $V$. For $r=1$, the $H_{1}$ term is given by the Schur functor associated to the cyclic Lie operad (cf. [Lev02], for example):
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V))\cong\mathbb{C}\mathrm{ycLie}(V).$
One can restrict to studying $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$, since the homological degree zero term can be derived from this. This is advantageous, as observed by Gadish and Hainaut, since $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ is usually smaller than $H_{0}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ in a precise sense.
The functor $V\mapsto H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ is analytic (in the sense of Joyal [Joy86]). In particular, for each $n\in\mathbb{N}$, one can focus on the $n$th homogeneous component, denoted here by $(-)^{[n]}$. The identification for $n\leq r+1$ is straightforward (see Proposition 3.2); there are $\mathfrak{S}_{r}$-equivariant natural isomorphisms
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}=\left\{\begin{array}[]{ll}0&n\leq r\\ \Gamma^{r+1}(V)\boxtimes\mathrm{triv}_{r}&n=r+1,\end{array}\right.$
where $\mathrm{triv}_{r}$ is the trivial representation of $\mathfrak{S}_{r}$ and $\Gamma^{r+1}(V)=(V^{\otimes r+1})^{\mathfrak{S}_{r+1}}$ is the $(r+1)$st divided power functor.
The purpose of this paper is to treat the next case, $n=r+2$. In the following statement, $\mathrm{sgn}_{r}$ denotes the sign representation of $\mathfrak{S}_{r}$, $\Lambda^{r+2}$ is the $(r+2)$nd exterior power functor and $\mathbf{S}_{(r,1^{2})}$ is the Schur functor associated to the partition $(r,1^{2})$.
Theorem 1.
For $r\in\mathbb{N}$, there are $\mathfrak{S}_{r}$-equivariant natural isomorphisms
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}\cong\left\{\begin{array}[]{ll}0&r=0\\ \Lambda^{3}(V)\boxtimes\mathrm{sgn}_{1}&r=1\\ \Lambda^{r+2}(V)\boxtimes\mathrm{sgn}_{r}\ \oplus\ \mathbf{S}_{(r,1^{2})}(V)\boxtimes\mathrm{triv}_{r}&r>1.\end{array}\right.$
Remark 1.
Using the Schur correspondence between representations of symmetric groups and polynomial functors, the result of Theorem 1 can be expressed for $r>1$ as stating that the Lie algebra homology is encoded by the representation of $\mathfrak{S}_{r+2}\times\mathfrak{S}_{r}$:
(1) $S_{(1^{r+2})}\boxtimes S_{(1^{r})}\ \oplus\ S_{(r,1^{2})}\boxtimes S_{(r)}.$
Theorem 1 is straightforward for $r\in\{0,1,2\}$. The case $r=3$ is established in Section 5. (This case was already known by computer calculation, for example as a consequence of calculations in [GH24]; it is established here by hand.) These cases serve as the initial steps for the inductive proof that is given in Section 6.
A consequence of Theorem 1 is:
Corollary 1.
Gadish and Hainaut’s conjecture [GH24, Conjecture 6.7] holds.
Let us outline how Corollary 1 follows from Theorem 1. Gadish and Hainaut’s conjecture is stated in terms of a bimodule that they denote $\Phi^{1}[r+2,r]$ (using the indexing $r$ compatible with that of Theorem 1). Their conjecture asserts that there is an isomorphism of $\mathfrak{S}_{r+2}\times\mathfrak{S}_{r}$-modules
(2) $\Phi^{1}[r+2,r]\cong S_{(r+2)}\boxtimes S_{(r)}\ \oplus\ S_{(3,1^{r-1})}\boxtimes S_{(1^{r})}$
(using this paper’s notation for the simple representations). Observe that the right hand side of (2) is obtained from (1) by replacing each occurrence of $S_{\lambda}$ for a partition $\lambda$ by $S_{\lambda^{\dagger}}$, where $\lambda^{\dagger}$ is the conjugate partition to $\lambda$.
In [GH24, Section 4.4], Gadish and Hainaut explain how their objects $\Phi[-,-]$ are related to higher Hochschild homology as considered in [PV25] and by Turchin and Willwacher in [TW19]. In particular, this explains how $\Phi^{1}[r+2,r]$ is related to $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$, also using the relationship between higher Hochschild homology and the Lie algebra homology considered here that is explained by the results of [PV25] in conjunction with those of [Pow24a] and [Pow24b]. In particular, as already observed in [PV25], this accounts for the twisting that results in the passage $\lambda\mapsto\lambda^{\dagger}$ observed above.
Acknowledgement
This project grew out of joint work with Christine Vespa; the author is very grateful for her ongoing interest. The author is also grateful to Louis Hainaut for his interest and his comments on a previous version.
2. Conventions and background
2.1. Representations of the symmetric groups and Schur functors
We work over the rational numbers, $\mathbb{Q}$. The category of finite-dimensional $\mathbb{Q}$-vector spaces is denoted by $\mathbf{mod}_{\mathbb{Q}}$; all undecorated tensor products $\otimes$ are understood to be taken over $\mathbb{Q}$. For $n\in\mathbb{N}$, the symmetric group on $n$ letters is denoted $\mathfrak{S}_{n}$. The category of $\mathbb{Q}\mathfrak{S}_{n}$-modules (or $\mathfrak{S}_{n}$-modules) is semi-simple. The simple $\mathfrak{S}_{n}$-modules are indexed by the partitions $\lambda$ of $n$, with $S_{\lambda}$ denoting the associated simple module. The indexing convention is chosen so that $S_{(n)}$ is the trivial representation $\mathrm{triv}_{n}$ and $S_{(1^{n})}$ the sign representation $\mathrm{sgn}_{n}$. For groups $H\subset G$, the notation $\downarrow^{G}_{H}$ denotes restriction from $G$-modules to $H$-modules and $\uparrow_{H}^{G}$ denotes induction.
The Schur functor construction establishes the correspondence between $\mathfrak{S}_{n}$-modules and functors from $\mathbf{mod}_{\mathbb{Q}}$ to $\mathbb{Q}$-vector spaces. For $M$ a $\mathfrak{S}_{n}$-module, the associated Schur functor is
$V\mapsto V^{\otimes n}\otimes_{\mathfrak{S}_{n}}M,$
for $V\in\mathbf{mod}_{\mathbb{Q}}$, where $\mathfrak{S}_{n}$ acts by place permutations on $V^{\otimes n}$. Writing $T^{n}$ for the $n$th tensor product functor $V\mapsto V^{\otimes n}$ equipped with the place permutation action of $\mathfrak{S}_{n}$ (on the right), the Schur functor can be written as
$T^{n}\otimes_{\mathfrak{S}_{n}}M.$
This is a polynomial functor that is homogeneous of degree $n$. (See [Mac15, Appendix I.A] for polynomial functors in characteristic zero.) The Schur functor construction establishes an equivalence of categories between $\mathfrak{S}_{n}$-modules and homogeneous polynomial functors of degree $n$; given such a polynomial functor $F$, the associated $\mathfrak{S}_{n}$-representation is given by the natural transformations
$\mathrm{Nat}(T^{n},F),$
where the $\mathfrak{S}_{n}$-action is given by the action on $T^{n}$. This can also be interpreted as a cross-effect functor, corresponding to passage to multilinear terms.
Notation 2.1.
For a partition $\lambda\vdash n$, the associated Schur functor is denoted $\mathbf{S}_{\lambda}$, so that $\mathbf{S}_{\lambda}(V)=V^{\otimes n}\otimes_{\mathfrak{S}_{n}}S_{\lambda}$.
Example 2.2.
For $\lambda=(n)$, the Schur functor $\mathbf{S}_{(n)}$ identifies as the $n$th symmetric power functor $V\mapsto S^{n}(V)=(V^{\otimes n})/\mathfrak{S}_{n}$. Over $\mathbb{Q}$, this is isomorphic to the $n$th divided power functor $V\mapsto\Gamma^{n}(V)=(V^{\otimes n})^{\mathfrak{S}_{n}}$. In particular, there is a canonical surjection $V^{\otimes n}\twoheadrightarrow S^{n}(V)$ and a canonical inclusion $\Gamma^{n}(V)\hookrightarrow V^{\otimes n}$. For $\lambda=(1^{n})$, the Schur functor $\mathbf{S}_{(1^{n})}$ identifies as the $n$th exterior power functor $V\mapsto\Lambda^{n}(V)$.
The $\mathbb{N}$-graded functors $S^{*}$, $\Gamma^{*}$ and $\Lambda^{*}$ are all exponential functors (paying attention to signs in the case of $\Lambda^{*}$). In particular, for $a,b\in\mathbb{N}$ and $n=a+b$, there are natural coproducts $\Lambda^{n}\rightarrow\Lambda^{a}\otimes\Lambda^{b}$ and $\Gamma^{n}\rightarrow\Gamma^{a}\otimes\Gamma^{b}$, and natural products $\Lambda^{a}\otimes\Lambda^{b}\rightarrow\Lambda^{n}$ and $S^{a}\otimes S^{b}\rightarrow S^{n}$.
We record the following basic fact which plays the rôle of Schur’s lemma:
Proposition 2.3.
For partitions $\lambda\vdash m$, $\mu\vdash n$,
  1. (1)
    $\mathbf{S}_{\lambda}$ is a simple functor from $\mathbf{mod}_{\mathbb{Q}}$ to $\mathbb{Q}$-vector spaces;
  2. (2)
    $\mathrm{Nat}(\mathbf{S}_{\lambda},\mathbf{S}_{\mu})\cong\left\{\begin{array}[]{ll}\mathbb{Q}&\lambda=\mu\\ 0&\mbox{otherwise.}\end{array}\right.$
For a partition $\lambda$, $l(\lambda)$ denotes its length, i.e., the number of non-zero terms in the partition. One has the following standard result on the connectivity of the simple Schur functors:
Lemma 2.4.
For a partition $\lambda$, $\mathbf{S}_{\lambda}(V)$ is zero if and only if $\dim V<l(\lambda)$.
We will also be working with functors from $\mathbf{mod}_{\mathbb{Q}}$ to $\mathbb{Q}$-vector spaces that are equipped with a $\mathfrak{S}_{r}$-action (for some fixed $r\in\mathbb{N}$) and $\mathfrak{S}_{r}$-equivariant natural transformations. A basic example of such a functor is given by forming the tensor product of a homogeneous polynomial functor $F$ of degree $n\in\mathbb{N}$ with a $\mathfrak{S}_{r}$-module $M$. This will be denoted by $F\boxtimes M$.
We also work with analytic functors (in the sense of [Joy86, Chapitre 4]). By definition, these are direct sums of homogeneous polynomial functors (not necessarily of the same degree).
Notation 2.5.
For $V\mapsto F(V)$ an analytic functor on $\mathbf{mod}_{\mathbb{Q}}$ and $n\in\mathbb{N}$, write $F^{[n]}$ for the homogeneous component of degree $n$, so that one has the natural isomorphism
$F(V)\cong\bigoplus_{n\in\mathbb{N}}F^{[n]}(V).$
Thus, written without the variable $V$, $F\cong\bigoplus_{n\in\mathbb{N}}F^{[n]}$. (We will also write $F(V)^{[n]}$ for $F^{[n]}(V)$.)
Example 2.6.
Consider the free Lie algebra functor $V\mapsto\mathbb{L}\mathrm{ie}(V)$, for $V\in\mathbf{mod}_{\mathbb{Q}}$. The underlying functor to $\mathbb{Q}$-vector spaces is analytic and one has the natural splitting
$\mathbb{L}\mathrm{ie}(V)\cong\bigoplus_{n\in\mathbb{N}}\mathbb{L}\mathrm{ie}^{[n]}(V).$
Here $\mathbb{L}\mathrm{ie}^{[n]}(V)$ identifies as the subspace of $\mathbb{L}\mathrm{ie}(V)$ generated (as a vector space) by iterated brackets of length exactly $n$. In particular, $\mathbb{L}\mathrm{ie}^{[0]}(V)=0$, $\mathbb{L}\mathrm{ie}^{[1]}(V)=V$, $\mathbb{L}\mathrm{ie}^{[2]}(V)\cong\Lambda^{2}(V)$ and $\mathbb{L}\mathrm{ie}^{[3]}(V)\cong\mathbf{S}_{(2,1)}(V)$.
2.2. A form of the de Rham differential
For later use, we introduce $\mathsf{d}$, a form of the de Rham differential written using the divided power functors. For $a,b\in\mathbb{N}$, with $b>0$, $\mathsf{d}$ is the natural transformation
(3) $\mathsf{d}\colon\Lambda^{a}\otimes\Gamma^{b}\rightarrow\Lambda^{a+1}\otimes\Gamma^{b-1}$
that is given by the composite
$\Lambda^{a}\otimes\Gamma^{b}\rightarrow\Lambda^{a}\otimes\Gamma^{1}\otimes\Gamma^{b-1}\cong\Lambda^{a}\otimes\Lambda^{1}\otimes\Gamma^{b-1}\rightarrow\Lambda^{a+1}\otimes\Gamma^{b-1}$
in which the first map is the coproduct of $\Gamma^{*}$ (which corresponds to the natural inclusion $(V^{\otimes b})^{\mathfrak{S}_{b}}\subset V\otimes(V^{\otimes b-1})^{\mathfrak{S}_{b-1}}$) and the second is the product of $\Lambda^{*}$.
One has the following characterization of $\mathsf{d}$:
Lemma 2.7.
For $a<b\in\mathbb{N}$, $\mathsf{d}$ is the unique map which fits into the commutative diagram of natural transformations
where the bottom map is induced by the product $\mu:\Lambda^{a}\otimes\Lambda^{1}\rightarrow\Lambda^{a+1}$ and the vertical maps correspond are induced by the canonical inclusion $\Gamma^{*}\subset T^{*}$.
As usual, one can identify the kernel and cokernel of $\mathsf{d}$ using the Pieri rule:
Lemma 2.8.
For positive integers $a,b$, the kernel and cokernel of $\mathsf{d}$ identify respectively as
$\ker(\mathsf{d})$$\cong$ $\mathbf{S}_{(b+1,1^{a-1})}$
$\mathrm{coker}(\mathsf{d})$$\cong$ $\mathbf{S}_{(b-1,1^{a+1})},$
where, if $b=1$, the cokernel is understood to be zero.
Example 2.9.
For $r\geq 2$, $\mathbf{S}_{(r,1^{2})}$ identifies as the kernel of $\mathsf{d}\colon\Lambda^{3}\otimes\Gamma^{r-1}\rightarrow\Lambda^{4}\otimes\Gamma^{r-2}.$
3. The adjoint action and a reduction
The main purpose of this section is to reduce the calculation of $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ (for $r\geq 2$) to that of the kernel of the morphism $\overline{\delta}_{r}$ introduced in Notation 3.7; this is achieved in Proposition 3.9.
3.1. The adjoint action
Consider $\mathbb{L}\mathrm{ie}(V)$, the free Lie algebra on $V\in\mathrm{Ob\ }\mathbf{mod}_{\mathbb{Q}}$. Then $\mathbb{L}\mathrm{ie}(V)$ can be considered as a left $\mathbb{L}\mathrm{ie}(V)$-module for the adjoint action, and this structure is natural with respect to $V$. Hence, for each $r\in\mathbb{N}$, $\mathbb{L}\mathrm{ie}(V)^{\otimes r}$ is a left $\mathbb{L}\mathrm{ie}(V)$-module (using the tensor product of $\mathbb{L}\mathrm{ie}(V)$-modules). For $r=0$ this gives the trivial module $\mathbb{Q}$ and, for $r=1$, the adjoint representation. The symmetric group $\mathfrak{S}_{r}$ acts by place permutations of the tensor factors of $\mathbb{L}\mathrm{ie}(V)^{\otimes r}$, respecting the $\mathbb{L}\mathrm{ie}(V)$-module structure. All this structure is natural with respect to $V$.
It follows that one can consider the Lie algebra homology: $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})$ and this takes values in $\mathfrak{S}_{r}$-modules, naturally with respect to $V$. This is given as the homology of the complex (placed in homological degrees $1$ and 00)
$V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r}\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{r}}}{{\longrightarrow}}\mathbb{L}\mathrm{ie}(V)^{\otimes r}$
given by restricting the adjoint action $\mathrm{ad}_{r}:\mathbb{L}\mathrm{ie}(V)\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r}\rightarrow\mathbb{L}\mathrm{ie}(V)^{\otimes r}$ to $V\subset\mathbb{L}\mathrm{ie}(V)$ on the first tensor factor. Thus one has
$H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})\cong\left\{\begin{array}[]{ll}\mathrm{coker}(\overline{\mathrm{ad}}_{r})&r=0\\ \ker(\overline{\mathrm{ad}}_{r})&r=1\\ 0&\mbox{otherwise}.\end{array}\right.$
Since $\overline{\mathrm{ad}}_{r}$ is a natural transformation between analytic functors (with respect to $V$), both $\ker(\overline{\mathrm{ad}}_{r})$ and $\mathrm{coker}(\overline{\mathrm{ad}}_{r})$ are analytic functors.
Example 3.1.
In the case $r=0$, the complex reduces to $V\stackrel{{\scriptstyle 0}}{{\rightarrow}}\mathbb{Q},$ giving the homology $H_{*}(\mathbb{L}\mathrm{ie}(V);\mathbb{Q})$ with trivial coefficients.
The case $r=1$ is more interesting; $\overline{\mathrm{ad}}_{1}$ fits into an exact complex:
$0\rightarrow\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{1}}}{{\rightarrow}}\mathbb{L}\mathrm{ie}(V)\rightarrow V\rightarrow 0$
Here, $\mathbb{C}\mathrm{ycLie}(V)$ is the Schur functor that is associated to the cyclic Lie operad. (Cf. [Lev02], where a stronger result is established, working over $\mathbb{Z}$).
To explain this exact sequence, recall that, if $\mathrm{Lie}(n)$ denotes the $n$th Lie representation (corresponding to the functor $V\mapsto\mathbb{L}\mathrm{ie}(V)^{[n]}$ by the Schur correspondence) and $\mathrm{CycLie}(n)$ the analogous module corresponding to $V\mapsto\mathbb{C}\mathrm{ycLie}(V)^{[n]}$, then these are related by
$\mathrm{Lie}(n)=\mathrm{CycLie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$
(this identification explains our conventions for indexing cyclic operads). The collection of morphisms $\mathrm{CycLie}(n+1)\rightarrow\mathrm{Lie}(n)\uparrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}$ obtained by adjunction induce the natural transformation $\mathbb{C}\mathrm{ycLie}(V)\rightarrow V\otimes\mathbb{L}\mathrm{ie}(V)$ above and it is straightforward to check that one obtains a sequence as given. Exactness can be checked at the level of representations of the symmetric groups by using restriction, exploiting the fact that $\mathrm{Lie}(n+1)\downarrow^{\mathfrak{S}_{n+1}}_{\mathfrak{S}_{n}}\cong\mathbb{Q}\mathfrak{S}_{n}$.
Here we are interested in $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}$ for small $n$, considered as a functor of $V\in\mathrm{Ob\ }\mathbf{mod}_{\mathbb{Q}}$ equipped with a $\mathfrak{S}_{r}$-action. Using the notation introduced in Section 2, one has the basic result:
Proposition 3.2.
For $r,n\in\mathbb{N}$, there is a $\mathfrak{S}_{r}$-equivariant natural isomorphism:
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}=\left\{\begin{array}[]{ll}0&n\leq r\\ \Gamma^{r+1}(V)\boxtimes\mathrm{triv}_{r}&n=r+1.\end{array}\right.$
Proof.
The case $n\leq r$ is clear, since $(V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[n]}$ is zero for $n\leq r$. The case $n=r+1$ is shown as follows. The case $r=0$ is immediate. For $r\geq 1$, one uses the the fact that (for $r\geq 1$) $\overline{\mathrm{ad}}_{r}^{[r+1]}$ identifies as an explicit map
(9) $V\otimes V^{\otimes r}\rightarrow\bigoplus_{i=0}^{r-1}V^{\otimes i}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-1-i}.$
The canonical inclusion $\Gamma^{r+1}(V)\hookrightarrow V\otimes V^{\otimes r}$ clearly maps to the kernel of this; moreover, the image is invariant under the action of $\mathfrak{S}_{r}$, by definition of the divided power functor. Hence, to prove the result, it suffices to show that, neglecting the $\mathfrak{S}_{r}$-action, the kernel is $\Gamma^{r+1}(V)$.
For $r=1$, the map (9) gives the natural short exact sequence
$0\rightarrow\Gamma^{2}(V)\rightarrow V^{\otimes 2}\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{1}}}{{\rightarrow}}\Lambda^{2}(V)\rightarrow 0,$
identifying the kernel as required. The cases $r>1$ then follow by an inductive argument (cf. the proof of Lemma 6.1). ∎
The case $n=r+2$ stated below is the main result of this paper (Theorem 1 of the Introduction).
Theorem 3.3.
For $r\in\mathbb{N}$, there is a $\mathfrak{S}_{r}$-equivariant natural isomorphism:
$H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}\cong\left\{\begin{array}[]{ll}0&r=0\\ \Lambda^{3}(V)\boxtimes\mathrm{sgn}_{1}&r=1\\ \Lambda^{r+2}(V)\boxtimes\mathrm{sgn}_{r}\ \oplus\ \mathbf{S}_{(r,1^{2})}(V)\boxtimes\mathrm{triv}_{r}&r>1.\end{array}\right.$
Example 3.4.
The case $r=1$ of Theorem 3.3 is elementary. The requisite complex is
$V\otimes\Lambda^{2}(V)\stackrel{{\scriptstyle\overline{\mathrm{ad}}_{1}^{[3]}}}{{\twoheadrightarrow}}\mathbf{S}_{(2,1)}(V),$
using the identifications $\mathbb{L}\mathrm{ie}(V)^{[2]}=\Lambda^{2}(V)$ and $\mathbb{L}\mathrm{ie}(V)^{[3]}=\mathbf{S}_{(2,1)}(V)$; the map is surjective since $\mathbb{L}\mathrm{ie}(V)$ is generated as a Lie algebra by $V$. By the Pieri rule, $V\otimes\Lambda^{2}(V)\cong\Lambda^{3}(V)\oplus\mathbf{S}_{(2,1)}(V)$, giving the identification of the kernel.
3.2. Reduction of the problem
To calculate $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$, we need to calculate the kernel of
$\overline{\mathrm{ad}}_{r}^{[r+2]}\ :\ \Big{(}V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r}\Big{)}^{[r+2]}\rightarrow\Big{(}\mathbb{L}\mathrm{ie}(V)^{\otimes r}\Big{)}^{[r+2]}.$
This can be described explicitly by using the identifications in Example 2.6.
Lemma 3.5.
For $r\geq 1$, there are natural $\mathfrak{S}_{r}$-equivariant isomorphisms
$\Big{(}V\otimes\mathbb{L}\mathrm{ie}(V)^{\otimes r}\Big{)}^{[r+2]}$$\cong$ $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
$\Big{(}\mathbb{L}\mathrm{ie}(V)^{\otimes r}\Big{)}^{[r+2]}$$\cong$ $(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\ \quad\oplus\ \quad(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}},$
for the obvious actions of the symmetric groups by place permutations (if $r=1$, the final term is understood to be zero).
The adjoint action $\overline{\mathrm{ad}}_{1}$ gives the canonical projections
[…]
with kernels $\Gamma^{2}(V)$ and $\Lambda^{3}(V)$ respectively. These extend to define the components $\delta^{\prime}_{r}$ and $\delta^{\prime\prime}_{r}$ of $\overline{\mathrm{ad}}_{r}^{[r+2]}$:
$\delta^{\prime}_{r}$ ::: $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$
$\delta^{\prime\prime}_{r}$ ::: $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}.$
Using the $\mathfrak{S}_{r}$-equivariance, these maps can be described by specifying one of their components, as follows.
The morphism underlying $\delta^{\prime}_{r}$ identifies as
$\bigoplus_{i=1}^{r}V\otimes(V^{\otimes i-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-i})\rightarrow\bigoplus_{j=1}^{r}(V^{\otimes j-1}\otimes\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-j}).$
The component with $i=j=1$ is given by
$V\otimes\Lambda^{2}(V)\otimes V^{\otimes r-1}\rightarrow\mathbf{S}_{(2,1)}(V)\otimes V^{\otimes r-1}$
obtained by applying $\overline{\mathrm{ad}}_{1}$ to the first two tensor factors of the domain.
Likewise, the morphism underlying $\delta^{\prime\prime}_{r}$ identifies as
$\bigoplus_{i=1}^{r}V\otimes(V^{\otimes i-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-i})\rightarrow\bigoplus_{1\leq j<k\leq r}(V^{\otimes j-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes k-j-1}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-k}).$
The component with $i=j=1$ and $k=2$ is
$V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\rightarrow\Lambda^{2}(V)\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2}$
obtained by applying $\overline{\mathrm{ad}}_{1}$ to the first and third tensor factors of the domain.
Lemma 3.6.
The kernel of $\delta^{\prime}_{r}$ is $\mathfrak{S}_{r}$-equivariantly naturally isomorphic to $(\Lambda^{3}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}},$ which is considered as a $\mathfrak{S}_{r}$-submodule of $V\otimes(\Lambda^{2}(V)\otimes V^{\otimes r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$ via the inclusion induced up from the $\mathfrak{S}_{r-1}$-equivariant map
$\Lambda^{3}(V)\otimes V^{\otimes r-1}\hookrightarrow V\otimes\Lambda^{2}V\otimes V^{\otimes r-1}$
given by applying the inclusion $\Lambda^{3}(V)\rightarrow V\otimes\Lambda^{2}(V)$ to the first factor.
Proof.
This generalizes the case $r=1$ given in Example 3.4. ∎
Notation 3.7.
Write $\overline{\delta}_{r}$ for the restriction of $\delta^{\prime\prime}_{r}$ to the kernel of $\delta^{\prime}_{r}$:
$\overline{\delta}_{r}\ :\ (\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}\rightarrow(\Lambda^{2}\otimes\Lambda^{2}\otimes T^{r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}.$
The underlying functors of the domain and codomain of $\overline{\delta}_{r}$ identify as follows:
Lemma 3.8.
There are isomorphisms of functors
$(\Lambda^{3}\otimes T^{r-1})\uparrow_{\mathfrak{S}_{r-1}}^{\mathfrak{S}_{r}}$$\cong$ $\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
$(\Lambda^{2}\otimes\Lambda^{2}\otimes T^{r-2})\uparrow_{\mathfrak{S}_{2}\times\mathfrak{S}_{r-2}}^{\mathfrak{S}_{r}}$$\cong$ $\bigoplus_{1\leq i<j\leq r}T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}.$
Thus, $\overline{\delta}_{r}$ identifies as a natural transformation
$\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\rightarrow\bigoplus_{1\leq i<j\leq r}T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}.$
Putting the above together provides the following reduction to the study of $\ker\overline{\delta}_{r}$.
Proposition 3.9.
For $r\geq 2$, $H_{1}(\mathbb{L}\mathrm{ie}(V);\mathbb{L}\mathrm{ie}(V)^{\otimes r})^{[r+2]}$ is isomorphic to $\ker\overline{\delta}_{r}$.
3.3. A geometric visualization
It is useful to visualize the maps in $\overline{\delta}_{r}$ using the $1$-skeleton of an $r-1$-simplex, considered as a directed bipartite graph as follows. The term $V^{\otimes i}\otimes\Lambda^{3}(V)\otimes V^{\otimes r-1-i}$ labels the black vertex of the $(r-1)$-simplex labelled by $i\in\{0,\ldots r-1\}$). For each edge of the $1$-skeleton, one adds a white vertex; on the edge between the black vertices $i\leq j$ this is labelled by the term $V^{\otimes i}\otimes\Lambda^{2}(V)\otimes V^{\otimes j-i}\otimes\Lambda^{2}(V)\otimes V^{\otimes r-2-j}$. This white vertex receives directed edges as indicated in the example below for $r=3$:
$\scriptstyle{\Lambda^{1}\otimes\Lambda^{2}\otimes\Lambda^{2}}$$\scriptstyle{\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{2}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}}$.$\scriptstyle{\Lambda^{1}\otimes\Lambda^{1}\otimes\Lambda^{3}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{1}\otimes\Lambda^{2}}$
This allows visualization of ‘restriction’ corresponding to faces. For instance, if one forgets all vertices that have either $\Lambda^{2}(V)$ or $\Lambda^{3}(V)$ in the last factor, one is left with:
$\scriptstyle{\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{2}\otimes\Lambda^{2}\otimes\Lambda^{1}}$$\scriptstyle{\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}}$.
This is obtained from the corresponding diagram for $r=2$ by applying $\otimes\Lambda^{1}$.
4. Analysing $\overline{\delta}_{r}$
This section first explains how to analyse the kernel of $\overline{\delta}_{r}$ (see Proposition 4.15). This is based on the morphism $\alpha:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}$ that is introduced in Section 4.3 (for the case $r=2$), as well as its analogues $\alpha_{i;j}$ for $r\geq 2$, where $1\leq i<j\leq r$ introduced in Section 4.4.
This is applied in Theorem 4.18 to establish the case $r=2$ of Theorem 1. It is also used in Proposition 4.20 to establish a lower bound for $\ker\overline{\delta}_{r}$ that will subsequently be shown to be an equality.
4.1. The functors $\Lambda^{3}\otimes\Lambda^{1}$ and $\Lambda^{2}\otimes\Lambda^{2}$
We start by reviewing the decomposition of $\Lambda^{3}\otimes\Lambda^{1}$ and $\Lambda^{2}\otimes\Lambda^{2}$ into simple functors, paying attention to the $\mathfrak{S}_{2}$-action in the second case.
Lemma 4.1.
There is a canonical splitting $\Lambda^{3}\otimes\Lambda^{1}\cong\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$ and a canonical $\mathfrak{S}_{2}$-equivariant splitting
$\Lambda^{2}\otimes\Lambda^{2}\cong\big{(}\Lambda^{4}\oplus\mathbf{S}_{(2,2)}\big{)}\boxtimes\mathrm{triv}_{2}\ \oplus\ \mathbf{S}_{(2,1^{2})}\boxtimes\mathrm{sgn}_{2},$
where $\mathfrak{S}_{2}$ acts by place permutations on the left hand side.
Proof.
The splitting $\Lambda^{3}\otimes\Lambda^{1}\cong\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$ and the underlying splitting $\Lambda^{2}\otimes\Lambda^{2}\cong\Lambda^{4}\oplus\mathbf{S}_{(2,2)}\oplus\mathbf{S}_{(2,1^{2})}$ follow from Pieri’s rule; they are canonical since the functors are multiplicity-free. It remains to consider how the $\mathfrak{S}_{2}$-action on $\Lambda^{2}\otimes\Lambda^{2}$ behaves with respect to the latter. By Schur’s lemma (Proposition 2.3 for Schur functors), this action corresponds to an invertible element of
(10) $\mathrm{End}(\Lambda^{4}\oplus\mathbf{S}_{(2,2)}\oplus\mathbf{S}_{(2,1^{2})})\cong\mathbb{Q}_{(1^{4})}\prod\mathbb{Q}_{(2,2)}\prod\mathbb{Q}_{(2,1^{2})},$
where the factors of the product of copies of $\mathbb{Q}$ are indexed by the partitions to which they correspond. More precisely, the element is of the form $(u,v,w)\in\mathbb{Q}\prod\mathbb{Q}\prod\mathbb{Q}$, where $u,v,w\in\{1,-1\}$.
There is an inclusion of $\Lambda^{4}$ into $\Lambda^{2}\otimes\Lambda^{2}$ given by the coproduct. The image is clearly invariant under transposition of tensor factors hence induces $\Lambda^{4}\hookrightarrow(\Lambda^{2}\otimes\Lambda^{2})^{\mathfrak{S}_{2}}$.
Since $(\Lambda^{2}\otimes\Lambda^{2})^{\mathfrak{S}_{2}}$ evaluated on $\mathbb{Q}^{\oplus 2}$ is clearly non-zero, it must also contain $\mathbf{S}_{(2,2)}$ as a composition factor, since $\mathbf{S}_{(2,1^{2})}$ and $\Lambda^{4}$ are both zero when evaluated on $\mathbb{Q}^{\oplus 2}$, by Lemma 2.4.
To conclude, it suffices to observe that $\Lambda^{2}\otimes\Lambda^{2}$ is not invariant under the $\mathfrak{S}_{2}$-action transposing factors. This implies that the remaining factor must correspond to $\mathbf{S}_{(2,1^{2})}\boxtimes\mathrm{sgn}_{2}$. (This can also be verified by an explicit calculation.) ∎
The direct sum decomposition of $\Lambda^{3}\otimes\Lambda^{1}$ can be made explicit as follows.
Lemma 4.2.
  1. (1)
    The idempotent corresponding to the projection of $\Lambda^{3}\otimes\Lambda^{1}$ onto the factor $\Lambda^{4}$ is
    $e:=\frac{1}{4}\Big{(}\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\mu}}{{\rightarrow}}\Lambda^{4}\stackrel{{\scriptstyle\psi}}{{\rightarrow}}\Lambda^{3}\otimes\Lambda^{1}\Big{)},$
    where $\mu$ is the product and $\psi$ the coproduct. Explicitly, $e(x\wedge y\wedge z\otimes w)=\frac{1}{4}\big{(}x\wedge y\wedge z\otimes w-x\wedge y\wedge w\otimes z+x\wedge z\wedge w\otimes y-y\wedge z\wedge w\otimes x\big{)}.$
  2. (2)
    The idempotent giving the projection onto $\mathbf{S}_{(2,1^{2})}$ is $(\mathrm{Id}-e)$. This is given explicitly by $(\mathrm{Id}-e)(x\wedge y\wedge z\otimes w)=\frac{1}{4}\big{(}3x\wedge y\wedge z\otimes w+x\wedge y\wedge w\otimes z-x\wedge z\wedge w\otimes y+y\wedge z\wedge w\otimes x\big{)}$.
Proof.
It is a straightforward verification that $e$ is an idempotent; since it factors across $\Lambda^{4}$ it gives the required projection. The second statement follows immediately. ∎
We record the following application of Schur’s lemma to endomorphisms of $\Lambda^{3}\otimes\Lambda^{1}$ analogous to 10 above:
Lemma 4.3.
The endomorphism ring of $\Lambda^{3}\otimes\Lambda^{1}\cong\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$ is canonically isomorphic to the product of two copies of $\mathbb{Q}$:
(11) $\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1})\cong\mathrm{End}(\Lambda^{4})\prod\mathrm{End}(\mathbf{S}_{(2,1^{2})})=\mathbb{Q}_{(1^{4})}\prod\mathbb{Q}_{(2,1^{2})}.$
4.2. The morphisms $\kappa$ and $\tilde{\kappa}$
We introduce the morphisms $\kappa$ and $\tilde{\kappa}$ that correspond to the components of $\overline{\delta}_{2}$:
Notation 4.4.
Let $\kappa:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$ denote the composite of the coproduct on the first factor followed by the product on the last factors:
(12) $\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\psi\otimes\mathrm{Id}}}{{\rightarrow}}\Lambda^{2}\otimes\Lambda^{1}\otimes\Lambda^{1}\stackrel{{\scriptstyle\mathrm{Id}\otimes\mu}}{{\rightarrow}}\Lambda^{2}\otimes\Lambda^{2}.$
The morphism $\tilde{\kappa}:\Lambda^{1}\otimes\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$ is defined by the commutative diagram:
${\Lambda^{1}\otimes\Lambda^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{\kappa}}$$\scriptstyle{\tau}$${\Lambda^{2}\otimes\Lambda^{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${\Lambda^{3}\otimes\Lambda^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa}$${\Lambda^{2}\otimes\Lambda^{2},}$
in which the maps labelled $\tau$ exchange the tensor factors.
The definition of $\tilde{\kappa}$ is justified by the following:
Lemma 4.5.
The natural transformation underlying $\overline{\delta}_{2}$ identifies as
$(\kappa,\tilde{\kappa})\ :\ \Lambda^{3}\otimes\Lambda^{1}\ \oplus\ \Lambda^{1}\otimes\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{2}.$
Proof.
The definition of $\tilde{\kappa}$ in terms of $\kappa$ ensures that $(\kappa,\tilde{\kappa})$ is $\mathfrak{S}_{2}$-equivariant. The result then follows from the characterization of $\overline{\delta}_{2}$. ∎
More generally, as for the description of $\delta_{r}^{\prime\prime}$ given above, one has:
Lemma 4.6.
For $r\geq 2$, the natural transformation $\overline{\delta}_{r}$ is the unique $\mathfrak{S}_{r}$-equivariant natural transformation with component
$\Lambda^{3}\otimes T^{r-1}\rightarrow\Lambda^{2}\otimes\Lambda^{2}\otimes T^{r-2}$
given by $\kappa\otimes\mathrm{Id}_{T^{r-2}}$.
4.3. The isomorphism $\alpha$
Before introducing $\alpha$, we require to establish some properties of $\kappa$ and $\tilde{\kappa}$:
Lemma 4.7.
The natural transformations
$\kappa$ ::: $\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
$\tilde{\kappa}$ ::: $\Lambda^{1}\otimes\Lambda^{3}\rightarrow\Lambda^{2}\otimes\Lambda^{2}$
are injective. Moreover $\mathrm{image}(\kappa)=\mathrm{image}(\tilde{\kappa})$.
Proof.
To prove that $\kappa$ is injective, using Lemma 4.1, it suffices to show that $\kappa|_{\Lambda^{4}}$ and $\kappa|_{\mathbf{S}_{(2,1^{2})}}$ are both non-trivial, since $\Lambda^{4}$ and $\mathbf{S}_{(2,1^{2})}$ are non-isomorphic simple functors (see Proposition 2.3). For $\Lambda^{4}$, it suffices to check that the composite $\Lambda^{4}\stackrel{{\scriptstyle\psi}}{{\rightarrow}}\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\kappa}}{{\rightarrow}}\Lambda^{2}\otimes\Lambda^{2}$ is non-trivial; this is a simple calculation. For $\mathbf{S}_{(2,1^{2})}$, it suffices to check that $\kappa$ is non-trivial when evaluated on $V=\mathbb{Q}^{\oplus 3}$; this is again straightforward.
That $\tilde{\kappa}$ is injective follows immediately from the definition of $\tilde{\kappa}$ in terms of $\kappa$. By inspection, again using Lemma 4.1, the images coincide. ∎
This gives the commutative diagram:
Definition 4.8.
Let $\alpha$ be the isomorphism $\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\cong}}{{\rightarrow}}\Lambda^{1}\otimes\Lambda^{3}$ induced by passage around the bottom of (17) (inverting the right hand isomorphism).
Thus $\tau\circ\alpha$ is an automorphism of $\Lambda^{3}\otimes\Lambda^{1}$, using the symmetry $\Lambda^{1}\otimes\Lambda^{3}\stackrel{{\scriptstyle\tau}}{{\rightarrow}}\Lambda^{3}\otimes\Lambda^{1}$. By Lemma 4.3, this identifies as follows:
Proposition 4.9.
Under the isomorphism (11), the automorphism $\tau\circ\alpha\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1})$ identifies as $(1,-1)$. In particular, $\tau\circ\alpha$ is an involution, i.e., $(\tau\circ\alpha)\circ(\tau\circ\alpha)=\mathrm{Id}$.
Proof.
Using the definition of $\tilde{\kappa}$ in terms of $\kappa$, this follows from Lemma 4.1, which describes the action of $\mathfrak{S}_{2}$ on $\Lambda^{2}\otimes\Lambda^{2}$ by place permutation in terms of the isotypical decomposition of $\Lambda^{2}\otimes\Lambda^{2}$. ∎
This allows the following explicit identification of the automorphism $\tau\circ\alpha$, using the idempotent $e\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1})$ of Lemma 4.2:
Corollary 4.10.
The automorphism $\tau\circ\alpha$ of $\Lambda^{3}\otimes\Lambda^{1}$ identifies as the morphism $2e-\mathrm{Id}$. Explicitly:
$\tau\circ\alpha(x\wedge y\wedge z\otimes w)=-\frac{1}{2}\big{(}x\wedge y\wedge z\otimes w+x\wedge y\wedge w\otimes z-x\wedge z\wedge w\otimes y+y\wedge z\wedge w\otimes x\big{)}.$
4.4. Extending $\alpha$ for $r>3$
The isomorphism $\alpha$ was introduced for the case $r=2$, where one only has to consider the terms $\Lambda^{3}\otimes\Lambda^{1}$ and $\Lambda^{1}\otimes\Lambda^{3}$. More generally, for each $1\leq i\leq r$, one has $T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ in which $\Lambda^{3}$ appears as the $i$th tensor factor.
Notation 4.11.
For integers $1\leq i<j\leq r$, let $\alpha_{i;j}$ denote the isomorphism
$\alpha_{i;j}:T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\rightarrow T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j}$
given by applying $\alpha:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}$ with respect to the $i$th and $j$th tensor factors and acting by the identity on the remaining factors.
Example 4.12.
For $r=2$, one has $\alpha_{1;2}=\alpha$. If $r>2$, then $\alpha_{1:2}$ is $\alpha\otimes\mathrm{Id}_{T^{r-2}}$.
Remark 4.13.
  1. (1)
    We only require to define $\alpha_{i;j}$ for $i<j$, since we can use $\alpha_{j;i}^{-1}$ to cover the cases $i>j$.
  2. (2)
    By Proposition 4.9, we have $\alpha_{j;i}^{-1}=\tau_{i;j}\circ\alpha_{j;i}\circ\tau_{i:j}$, where $\tau_{i;j}$ transposes the $i$th and $j$th tensor factors.
4.5. Identifying $\ker\overline{\delta}_{r}$
Suppose that $r\geq 2$. Motivated by Proposition 3.9, we seek to calculate $\ker\overline{\delta}_{r}$, where $\overline{\delta}_{r}$ is the morphism introduced in Notation 3.7. The latter has the form given in Lemma 3.8:
$\overline{\delta}_{r}\colon\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\rightarrow\bigoplus_{1\leq i<j\leq r}T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}.$
This is described explicitly by Lemma 4.6 in terms of $\kappa$.
Notation 4.14.
Write an element of $\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ as $(X_{i})_{1\leq i\leq r}$, where $X_{i}\in T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$.
Proposition 4.15.
Suppose that $r\geq 2$. An element $(X_{i})_{1\leq i\leq r}$ of $\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ lies in $\ker\overline{\delta}_{r}$ if and only if, for all $1\leq i<j\leq r$, one has
$\alpha_{i;j}(X_{i})=-X_{j}.$
Proof.
By Lemma 4.6, the definition of the $\alpha_{i,j}$ ensures that this result holds. Namely, by definition of $\overline{\delta}_{r}$, $(X_{i})_{1\leq i\leq r}$ lies in the kernel of $\overline{\delta}_{r}$ if and only if, for each $1\leq i<j\leq r$, the map
$(T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i})\ \oplus\ (T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j})\rightarrow T^{i-1}\otimes\Lambda^{2}\otimes T^{j-i-1}\otimes\Lambda^{2}\otimes T^{r-j}$
induced by $\kappa$ and $\tilde{\kappa}$ applied with respect to the $i$th and $j$th tensor factors is zero. Here $X_{i}$ lies in $T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$ and $X_{j}$ in $T^{j-1}\otimes\Lambda^{3}\otimes T^{r-j}$. The latter condition is equivalent to $\alpha_{i;j}(X_{i})=-X_{j}$, by definition of $\alpha_{i;j}$. ∎
This has the immediate consequence:
Corollary 4.16.
Suppose that $r\geq 2$. For any $1\leq\ell\leq r$, the projection
$\bigoplus_{i=1}^{r}T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}\twoheadrightarrow T^{\ell-1}\otimes\Lambda^{3}\otimes T^{r-\ell}$
onto the $\ell$th factor induces an inclusion
(18) $\ker\overline{\delta}_{r}\hookrightarrow T^{\ell-1}\otimes\Lambda^{3}\otimes T^{r-\ell}.$
Remark 4.17.
In general (as follows from Theorem 1, for example) the inclusion (18) is not an isomorphism. For instance, for $r=3$, we can choose any $X_{1}\in\Lambda^{3}\otimes T^{2}$ and then set $X_{2}:=-\alpha_{1;2}(X_{1})$ and $X_{3}:=-\alpha_{1;3}(X_{1})$. This corresponds to an element of $\ker\overline{\delta}_{3}$ if and only if $\alpha_{2;3}(X_{2})=-X_{3}.$ This condition is not satisfied for every $X_{1}$. The additional condition can be considered as a ‘cocycle’ condition.
However, for $r=2$, there is no ‘cocycle’ condition, so that the inclusion (18) gives an isomorphism $\ker\overline{\delta}_{2}\cong\Lambda^{3}\otimes\Lambda^{1}$, taking $\ell=1$. This can be strengthened to take into account the $\mathfrak{S}_{2}$-equivariance:
Theorem 4.18.
For $r=2$, there is a $\mathfrak{S}_{2}$-equivariant isomorphism:
$\ker\overline{\delta}_{2}\cong\Lambda^{4}\boxtimes\mathrm{sgn}_{2}\ \oplus\ \mathbf{S}_{(2,1^{2})}\boxtimes\mathrm{triv}_{2}.$
Moreover, up to non-zero scalar
  1. (1)
    the underlying inclusion $\Lambda^{4}\hookrightarrow\Lambda^{3}\otimes\Lambda^{1}\oplus\Lambda^{1}\otimes\Lambda^{3}$ is given by the coproduct of $\Lambda^{*}$;
  2. (2)
    the underlying inclusion $\mathbf{S}_{(2,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Lambda^{1}\oplus\Lambda^{1}\otimes\Lambda^{3}$ corresponds to a commutative diagram:
    ${\mathbf{S}_{(2,1^{2})}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Lambda^{3}\otimes\Lambda^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau}$${\Lambda^{1}\otimes\Lambda^{3}}$
    associated to $\mathbf{S}_{(2,1^{2})}\subset\Lambda^{3}\otimes\Lambda^{1}$.
Proof.
By the discussion preceding the statement and Pieri’s rule, we already know that the underlying functor of $\ker\overline{\delta}_{2}$ is $\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$. It remains to consider the $\mathfrak{S}_{2}$-action. This is analysed by using Proposition 4.9, which shows that $\tau\circ\alpha$ corresponds to the action of the diagonal matrix with entries $(1,-1)$. The explicit description of the underlying inclusions follows directly from this. ∎
Remark 4.19.
The sign in the equality $\alpha(X_{1})=-X_{2}$ appears in different ways for the cases $\Lambda^{4}$ and $\mathbf{S}_{(2,1^{2})}$. For $\Lambda^{4}$, it arises from the ‘graded commutativity’ of the coproduct of $\Lambda^{*}$. For $\mathbf{S}_{(2,1^{2})}$, the sign arises from $\tau\circ\alpha$ whereas the embedding of $\mathbf{S}_{(2,1^{2})}$ in $\Lambda^{1}\otimes\Lambda^{3}$ is given by the commutative triangle in the statement of the Theorem, using $\tau$.
4.6. A lower bound for $\ker\overline{\delta}_{r}$
Generalizing the case $r=2$, we establish a lower bound for $\ker\overline{\delta}_{r}$ for $r\geq 2$. This will later be shown to coincide with $\ker\overline{\delta}_{r}$.
Proposition 4.20.
For $r\geq 2$, there is a $\mathfrak{S}_{r}$-equivariant inclusion:
$\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}\ \oplus\ \mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}\hookrightarrow\ker\overline{\delta}_{r}.$
Proof.
The terms $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}$ and $\mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}$ are simple and non-isomorphic, hence it suffices to exhibit the appropriate inclusions for the individual factors.
For the underlying functor $\Lambda^{r+2}$, for each $i\in\{1,\ldots,r\}$, the natural transformation
$\Lambda^{r+2}\rightarrow T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
is given by the coproduct of $\Lambda^{*}$. Generalizing the case $r=2$ given in Theorem 4.18, one checks using coassociativity of the coproduct of $\Lambda^{*}$ that these induce an inclusion $\Lambda^{r+2}\subset\ker\overline{\delta}_{r}$. Moreover, taking into account the $\mathfrak{S}_{r}$-action, one checks that this defines a $\mathfrak{S}_{r}$-equivariant inclusion $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}\subset\ker\overline{\delta}_{r}$.
Similarly, for $\mathbf{S}_{(r,1^{2})}$, one starts from a chosen inclusion (unique up to non-zero scalar by Schur’s Lemma) $\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{r-1}$. Observe that, by Example 2.9 and the characterization of $\mathsf{d}$ given in Lemma 2.7, this fits into a commutative diagram
where the right hand vertical map is induced by $\mathbf{S}_{(2,1^{2})}\subset\Lambda^{3}\otimes\Gamma^{1}$ and the lower horizontal map by the coproduct of $\Gamma^{*}$.
For $1\leq i\leq r$, using the coproduct of $\Gamma^{*}$, one has the composite inclusion:
$\mathbf{S}_{(r,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{r-1}\rightarrow\Lambda^{3}\otimes\Gamma^{i-1}\otimes\Gamma^{r-i}\cong\Gamma^{i-1}\otimes\Lambda^{3}\otimes\Gamma^{r-i}\subset T^{i-1}\otimes\Lambda^{3}\otimes T^{r-i}$
where the isomorphism is given by the symmetric monoidal structure and the inclusion is given by $\Gamma^{*}\subset T^{*}$. Using the commutative diagram (23) and the coassociativity of the coproduct of $\Gamma^{*}$, one deduces from Theorem 4.18 that these maps define an inclusion $\mathbf{S}_{(r,1^{2})}\subset\ker\overline{\delta}_{r}$. Moreover, from the construction, it is clear that this corresponds to a $\mathfrak{S}_{r}$-equivariant inclusion $\mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}\subset\ker\overline{\delta}_{r}$. ∎
5. The case $r=3$
A key step in proving Theorem 1 is the case $r=3$:
Theorem 5.1.
There is a $\mathfrak{S}_{3}$-equivariant natural isomorphism
$\ker\overline{\delta}_{3}\cong\Lambda^{5}\boxtimes\mathrm{sgn}_{3}\ \oplus\ \mathbf{S}_{(3,1^{2})}\boxtimes\mathrm{triv}_{3}.$
The proof uses Proposition 4.15 for the case $r=3$. The relevant morphisms (see Notation 4.11) fit into the (non-commutative) diagram:
${\Lambda^{1}\otimes\Lambda^{1}\otimes\Lambda^{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{1;3}^{-1}}$${\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{1;2}}$${\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}.\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{2;3}}$
The cocycle condition arising from Proposition 4.15 can be restated as
Proposition 5.2.
The kernel of $\overline{\delta}_{3}$ is isomorphic to the equalizer of
${\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Id}}$$\scriptstyle{-(\alpha_{1;3}^{-1}\circ\alpha_{2;3}\circ\alpha_{1;2})}$   ${\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}.}$
The composite $\alpha_{1;3}^{-1}\circ\alpha_{2;3}\circ\alpha_{1;2}$ is an automorphism of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$. This will be analysed using:
Notation 5.3.
Denote by $\beta$ the automorphism of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ given by the composite
$\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\stackrel{{\scriptstyle\tau\circ\alpha\otimes\mathrm{Id}}}{{\longrightarrow}}\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\stackrel{{\scriptstyle\mathrm{Id}\otimes\tau}}{{\longrightarrow}}\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}.$
Remark 5.4.
The morphism $\beta$ also identifies as the composite of $\alpha\otimes\mathrm{Id}:\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}$ with the cyclic place permutation $\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\mathrm{cyclic}}}{{\rightarrow}}\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ (this corresponds to $\Lambda^{1}\otimes F\rightarrow F\otimes\Lambda^{1}$, where $F:=\Lambda^{3}\otimes\Lambda^{1}$.)
Lemma 5.5.
The composite $\alpha_{1;3}^{-1}\circ\alpha_{2;3}\circ\alpha_{1;2}$ is equal to the automorphism $\beta^{\circ 3}$ of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$.
Proof.
This is a direct verification, based on the construction of the morphisms $\alpha_{1;3}$, $\alpha_{2;3}$, $\alpha_{1;2}$ from $\alpha$. Namely, using Remark 5.4, one reduces to checking that the cyclic permutation in the definition of $\beta$ always ensures that the $\alpha$ is applied to the correct choice of factors, in the correct order. ∎
Using Lemma 5.5, Proposition 5.2 can be restated as:
Proposition 5.6.
The kernel of $\overline{\delta}_{3}$ is equal to the equalizer of $\mathrm{Id}$ and $(-\beta)^{\circ 3}$.
We thus need to understand the automorphism $-\beta$ of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$. This identifies explicitly using Corollary 4.10:
Lemma 5.7.
The automorphism $-\beta$ of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ evaluated on $V$ is given for $x,y,z,s,t\in V$ by
$-\beta(x\wedge y\wedge z\otimes s\otimes t)=\frac{1}{2}\big{(}x\wedge y\wedge z\otimes t\otimes s+x\wedge y\wedge s\otimes t\otimes z-x\wedge z\wedge s\otimes t\otimes y+y\wedge z\wedge s\otimes t\otimes x\big{)}.$
To analyse the automorphism $-\beta$ by exploiting the representation theory, one can use the natural decomposition of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ that is given by Pieri’s rule:
(24) $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\cong\Lambda^{5}\oplus\mathbf{S}_{(3,1^{2})}\oplus\mathbf{S}_{(2^{2},1)}\oplus\mathbf{S}_{(2,1^{3})}^{\oplus 2}.$
By Schur’s lemma, we have the following analogue of Lemma 4.3:
Lemma 5.8.
There is an isomorphism of algebras
$\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$$\cong$ $\mathrm{End}(\Lambda^{5})\prod\mathrm{End}(\mathbf{S}_{(3,1^{2})})\prod\mathrm{End}(\mathbf{S}_{(2^{2},1)})\prod\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})$
$\cong$ $\mathbb{Q}_{(1^{5})}\prod\mathbb{Q}_{(3,1^{2})}\prod\mathbb{Q}_{(2^{2},1)}\prod M_{2}(\mathbb{Q}).$
Remark 5.9.
The fact that the decomposition (24) is not multiplicity-free is the source of complexity in the proof of Theorem 5.1. In particular, the isomorphism $\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ is not canonical.
We analyse the isomorphism (24) in two different ways, adapted to considering the action of $\tau\circ\alpha\otimes\mathrm{Id}$ and of $\mathrm{Id}\otimes\tau$, the two isomorphisms appearing in the construction of $\beta$ in Notation 5.3.
For the first, starting from the decomposition $\Lambda^{3}\otimes\Lambda^{1}\cong\Lambda^{4}\oplus\mathbf{S}_{(2,1^{2})}$, one uses the identifications given by the Pieri rule
$\Lambda^{4}\otimes\Lambda^{1}$$\cong$ $\Lambda^{5}\oplus\mathbf{S}_{(2,1^{3})}$
$\mathbf{S}_{(2,1^{2})}\otimes\Lambda^{1}$$\cong$ $\mathbf{S}_{(3,1^{2})}\oplus\mathbf{S}_{(2^{2},1)}\oplus\mathbf{S}_{(2,1^{3})}.$
Then, by Corollary 4.9, one has the following, in which $(1,-1)$ denotes the diagonal matrix in $M_{2}(\mathbb{Q})$ with respect to the chosen decomposition:
Lemma 5.10.
There exists an isomorphism as in (24) with respect to which $\tau\circ\alpha\otimes\mathrm{Id}\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$ acts via the isomorphism of Lemma 5.8 by $(1,-1,-1,(1,-1))$.
To study the action of $\mathrm{Id}\otimes\tau$, one uses the $\mathfrak{S}_{2}$-equivariant decomposition $\Lambda^{1}\otimes\Lambda^{1}\cong\mathbf{S}_{(2)}\boxtimes\mathrm{triv}_{2}\oplus\Lambda^{2}\boxtimes\mathrm{sgn}_{2}$ (for the place permutation action on the left hand side) together with the isomorphisms given by the Pieri rule:
$\Lambda^{3}\otimes\mathbf{S}_{(2)}$$\cong$ $\mathbf{S}_{(3,1^{2})}\oplus\mathbf{S}_{(2,1^{3})}$
$\Lambda^{3}\otimes\Lambda^{2}$$\cong$ $\mathbf{S}_{(2^{2},1)}\oplus\mathbf{S}_{(2,1^{3})}\oplus\Lambda^{5}.$
Then, analogously to Lemma 5.10, one has:
Lemma 5.11.
There exists an isomorphism as in (24) with respect to which $\mathrm{Id}\otimes\tau\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$ acts via the isomorphism of Lemma 5.8 by $(-1,1,-1,(1,-1))$.
Since there are unique composition factors of $\Lambda^{5}$, $\mathbf{S}_{(3,1^{2})}$, and $\mathbf{S}_{(2^{2},1)}$ in $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$, Lemmas 5.10 and 5.11 imply:
Lemma 5.12.
The automorphism $-\beta$ restricts to
  1. (1)
    $\mathrm{Id}$ on the summand $\Lambda^{5}\oplus\mathbf{S}_{(3,1^{2})}$;
  2. (2)
    $-\mathrm{Id}$ on the summand $\mathbf{S}_{(2^{2},1)}$.
To prove Theorem 5.1, it remains to show that $(-\beta)^{\circ 3}$ has no non-zero fixed points for the factor $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$; this restriction can be considered as an element of $\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ via Lemma 5.8. It is sufficient to calculate the eigenvalues of this action.
Now, the decompositions corresponding to the two composition factors $\mathbf{S}_{(2,1^{3})}$ in $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ that are used in Lemmas 5.10 and 5.11 do not coincide (that they must be different is a consequence of the calculations below), so one cannot directly read off the action of $-\beta$ on the corresponding isotypical component. One can, however, deduce the determinant:
Lemma 5.13.
The automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has determinant $1$.
Proof.
Using Lemmas 5.10 and 5.11, one sees that the determinant of $\beta$ is $1$, using that the diagonal matrix $\mathrm{diag}(1,-1)$ has determinant $-1$. Thus $-\beta$ also has determinant $1$, since $-\mathrm{Id}=\mathrm{diag}(-1,-1)$ has determinant one. ∎
To determine the eigenvalues, it thus suffices to calculate the trace. For this, it is convenient to pass to the associated representations of $\mathfrak{S}_{5}$ and their underlying vector spaces, using the Schur correspondence that is given by the Schur functor construction, with inverse given by cross effects.
The Schur functor $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ corresponds to the $\mathfrak{S}_{5}$-representation
(25) $(\mathrm{sgn}_{3}\boxtimes\mathrm{triv}_{1}\boxtimes\mathrm{triv}_{1})\uparrow_{\mathfrak{S}_{3}\times\mathfrak{S}_{1}\times\mathfrak{S}_{1}}^{\mathfrak{S}_{5}}\cong\mathrm{sgn}_{5}\oplus S_{(3,1^{2})}\oplus S_{(2^{2},1)}\oplus S_{(2,1^{3})}^{\oplus 2},$
This module has dimension $20$; the dimensions of $\mathrm{sgn}_{5}$, $S_{(3,1^{2})}$, $S_{(2^{2},1)}$, and $S_{(2,1^{3})}$ are respectively $1$, $6$, $5$, and $4$. The automorphism $-\beta$ induces an automorphism of the $\mathfrak{S}_{5}$-module given by (25), and hence of its underlying $\mathbb{Q}$-vector space, $\mathbb{Q}^{\oplus 20}$.
Lemma 5.14.
The element of $M_{20}(\mathbb{Q})$ associated to $-\beta$ has trace 00.
Proof.
The vector space underlying (25) has basis given by the elements $(i\wedge j\wedge k)\otimes s\otimes t$, where $i,j,k,s,t$ are pairwise distinct elements of $\{1,\ldots,5\}$ with $i<j<k$. (This basis can be ordered lexicographically.)
By inspection using Lemma 5.7 and its analogue for the $\mathfrak{S}_{5}$-representations (given by considering cross-effects), the associated matrix is zero on the diagonal, hence the trace is zero. ∎
From this, one deduces the following counterpart of Lemma 5.13
Lemma 5.15.
The automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has trace $-\frac{1}{2}$.
Proof.
We first calculate the trace of the restriction of $-\beta$ (considered as an element of $M_{20}(\mathbb{Q})$ as in Lemma 5.14) to the direct summand $S_{(2,1^{3})}^{\oplus 2}$ (which has underlying vector space of dimension $8$). Using the decomposition (25) and Schur’s lemma, $-\beta$ has block diagonal form. Moreover, by Lemma 5.12, it acts as $\mathrm{Id}$ on $\mathrm{sgn}_{5}\oplus S_{(3,1^{2})}$ (which has dimension $7$) and by $-\mathrm{Id}$ on $S_{(2^{2},1)}$ (which has dimension $5$). Using the block diagonal form and Lemma 5.14, it follows that the restriction of $-\beta$ to $S_{(2,1^{3})}^{\oplus 2}$ (an element of $M_{8}(\mathbb{Q})$) has trace $-2$.
By Schur’s Lemma, this element of $M_{8}(\mathbb{Q})$ is obtained from $\mathrm{End}_{\mathfrak{S}_{5}}(S_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ by the morphism of algebras $M_{2}(\mathbb{Q})\rightarrow M_{8}(\mathbb{Q})$ induced by the functor $-\otimes\mathbb{Q}^{\oplus 4}$ on vector spaces. (The trace is independent of choices made.) The result follows. ∎
Proof of Theorem 5.1.
By Proposition 5.6, $\ker\overline{\delta}_{3}$ is the equalizer of $\mathrm{Id}$ and $(-\beta)^{\circ 3}$. The behaviour of this on the single-multiplicity composition factors is given by Lemma 5.12. To conclude, we need to show that $\mathrm{Id}-(-\beta)^{\circ 3}$ is injective restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$.
Using the trace and determinant, from Lemmas 5.13 and 5.15, one deduces that the automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has eigenvalues (in $\mathbb{C}$) $\frac{1}{4}(-1\pm\sqrt{15}i).$ This implies that $1$ is not an eigenvalue of $(-\beta)^{\circ 3}$, whence the result. ∎
6. The proof of Theorem 1
The purpose of this section is to prove Theorem 1, based on the analysis of $\ker\overline{\delta}_{r}$ (for $r\geq 2$) by applying Proposition 3.9. The proof will exploit the lemmas concerning Schur functors given in Section 6.1.
6.1. Preparatory lemmas
Consider the exterior power functors $\Lambda^{n}\cong\mathbf{S}_{(1^{n})}$, for $n\in\mathbb{N}$. The coproduct provides the inclusion $\Lambda^{n+1}\hookrightarrow\Lambda^{n}\otimes\Lambda^{1}$; up to non-zero scalar multiple, this is the unique non-trivial natural transformation of this form. On applying $-\otimes\Lambda^{1}$, one therefore obtains $\Lambda^{n+1}\otimes\Lambda^{1}\hookrightarrow\Lambda^{n}\otimes T^{2}$. Post-composing with $\mathrm{Id}_{\Lambda^{n}}\otimes\tau$ gives a second, distinct embedding.
Up to non-zero scalar multiple, there is a unique non-trivial map from $\Lambda^{n+2}$ to $\Lambda^{n}\otimes T^{2}$, so that $\Lambda^{n+2}$ is canonically a subfunctor of the codomain. These subobjects of $\Lambda^{n}\otimes T^{2}$ are related by:
Lemma 6.1.
For $n\geq 1$, the intersection of $\Lambda^{n+1}\otimes\Lambda^{1}$ and $(\mathrm{Id}\otimes\tau)\big{(}\Lambda^{n+1}\otimes\Lambda^{1}\big{)}$ in $\Lambda^{n}\otimes T^{2}$ is $\Lambda^{n+2}$.
Proof.
Write $\mathfrak{S}_{n+1}\subset\mathfrak{S}_{n+2}$ for the subgroup fixing $n+2$ and $\mathfrak{S}_{n+1}^{\prime}\subset\mathfrak{S}_{n+2}$ for that fixing $n+1$. Considering $\Lambda^{n+1}\otimes\Lambda^{1}$ and $(\mathrm{Id}\otimes\tau)\big{(}\Lambda^{n+1}\otimes\Lambda^{1}\big{)}$ as subobjects of $T^{n+2}$, the former identifies as the subspace of elements on which $\sigma\in\mathfrak{S}_{n+1}$ acts by multiplication by $\mathrm{sgn}(\sigma)$ and the latter similarly, with respect to $\mathfrak{S}_{n+1}^{\prime}$. Since the subgroups $\mathfrak{S}_{n+1}$ and $\mathfrak{S}_{n+1}^{\prime}$ generate $\mathfrak{S}_{n+2}$, it follows that any $\rho\in\mathfrak{S}_{n+2}$ acts on the intersection by multiplication by $\mathrm{sgn}(\rho)$. The result follows. ∎
We now consider the family of functors $\mathbf{S}_{(n,1^{2})}$, assuming that $n\geq 2$. One has the embedding
$\mathbf{S}_{(n,1^{2})}\hookrightarrow\Lambda^{3}\otimes\Gamma^{n-1}\hookrightarrow\Lambda^{3}\otimes\Gamma^{n-2}\otimes\Gamma^{1},$
where the first map is given by Example 2.9 and the second map by the coproduct of $\Gamma^{*}$.
Forming the tensor product $\otimes\Gamma^{1}$, one obtains the inclusion
(26) $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\subset\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}.$
Post-composing with $(\mathrm{Id}\otimes\tau)$ gives the inclusion $(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1})\subset\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$. The following result is a counterpart of Lemma 6.1:
Lemma 6.2.
For $n\geq 2$, the intersection of $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}$ and $(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\big{)}$ in $\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$ is isomorphic to $\mathbf{S}_{(n+1,1^{2})}$.
Proof.
One first observes that the intersection, denoted here simply by $\bigcap$, lies in $\Lambda^{3}(V)\otimes\Gamma^{n}(V)$. This follows from an argument similar to that used in the proof of Lemma 6.1.
It then remains to show (using the identification of $\mathbf{S}_{(n+1,1^{2})}$ given in Example 2.9) that this intersection lies in the kernel of $\mathsf{d}$.
By construction, there is a commutative diagram:
[…]
in which $\mu:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{4}$ is the product and the top square is given by the definition of the intersection $\cap$ and the above, using the embedding (26) composed with $\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}\subset\Lambda^{3}\otimes T^{n}$.
The labelled map is zero by Lemma 2.8, using the characterization of $\mathsf{d}$ given by Lemma 2.7. This implies that $\bigcap$ lies in the kernel of the right hand vertical composite. Since the latter factors across $\mathsf{d}\colon\Lambda^{3}\otimes\Gamma^{n}\rightarrow\Lambda^{4}\otimes\Gamma^{n-1}$, it follows that $\bigcap$ lies in $\ker\mathsf{d}\cong\mathbf{S}_{(n+1,1^{2})}$ (using Example 2.9). To conclude, it remains to show that this is an equality (this is equivalent to establishing the non-triviality of $\bigcap$).
By Pieri’s rule, $\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$ contains a unique composition factor of $\mathbf{S}_{(n+1,1^{2})}$; this occurs in both $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}$ and $(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\big{)}$, hence occurs in the intersection, thus concluding the proof. ∎
6.2. The proof of Theorem 1
The following proof exploits the description of $\ker\overline{\delta}_{r}$ given in Proposition 4.15.
Proof.
The cases $r\in\{1,2,3\}$ have been established in Example 3.4, Theorem 4.18, and Theorem 5.1 respectively. Hence, by induction upon $r$, we may suppose that $r>3$ and that the result has been proved for $r-1$. It remains to prove the inductive step.
Proposition 4.20 established the $\mathfrak{S}_{r}$-equivariant natural inclusion $\Lambda^{r+2}\boxtimes\mathrm{sgn}_{r}\ \oplus\ \mathbf{S}_{(r,1^{2})}\boxtimes\mathrm{triv}_{r}\subset\ker(\overline{\delta}_{r})$, hence it suffices only to consider the underlying functor, neglecting the $\mathfrak{S}_{r}$-action.
Based upon Proposition 4.15, Corollary 4.16 with $\ell=1$ gives the inclusion
$\ker\overline{\delta}_{r}\subset\Lambda^{3}\otimes T^{r-1}.$
Using Proposition 4.15, we claim that there are inclusions:
(27) $\ker\overline{\delta}_{r}$$\subset$ $\ker\overline{\delta}_{r-1}\otimes\Lambda^{1}$
(28) $\ker\overline{\delta}_{r}$$\subset$ $(\mathrm{Id}_{\Lambda^{3}\otimes T^{\otimes r-3}}\otimes\tau)\big{(}\ker\overline{\delta}_{r-1}\otimes\Lambda^{1}\big{)},$
where both terms are considered as subobjects of $\Lambda^{3}\otimes T^{r-1}$ and $\ker\overline{\delta}_{r-1}$ is considered as a subobject of $\Lambda^{3}\otimes T^{r-2}$, again by Corollary 4.16 with $\ell=1$. The $\tau$ corresponds to the action of $\mathfrak{S}_{2}$ by place permutations on $T^{2}$, the last two tensor factors of $T^{r-1}$.
To establish this claim, it is useful to visualize the statement of Proposition 4.15 by considering that $i\in\{1,\ldots,r\}$ indexes the vertices of a $r-1$-simplex, analogously to the geometric visualization proposed in Section 3.3. Then the inclusion (27) is obtained by restricting to the codimension one face given by $i\neq r$, using Proposition 4.15 for $r-1$ to identify $\ker\overline{\delta}_{r-1}$. The inclusion (3.3) is obtained similarly, this time using the face $i\neq r-1$. Using the equivariance property of $\overline{\delta}_{r}$, one obtains the stated inclusion by exploiting the transposition of tensor factors $\tau$.
Now, by the inductive hypothesis, there is a natural isomorphism:
$\ker\overline{\delta}_{r-1}\otimes\Lambda^{1}\cong\Lambda^{r+1}\otimes\Lambda^{1}\ \oplus\ \mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}.$
Hence, putting the information from the two faces (27) and (28) together gives
(29) $\ker\overline{\delta}_{r}\subset\big{(}\Lambda^{r+1}\otimes\Lambda^{1}\ \oplus\ \mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}\big{)}\cap(\mathrm{Id}\otimes\tau)\big{(}\Lambda^{r+1}\otimes\Lambda^{1}\ \oplus\ \mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}\big{)}$
where the intersection is formed as subobjects of $\Lambda^{3}\otimes T^{r-1}$.
The Pieri rule gives the natural isomorphisms:
$\Lambda^{r+1}\otimes\Lambda^{1}$$\cong$ $\Lambda^{r+2}\oplus\mathbf{S}_{(2,1^{r})}$
$\mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}$$\cong$ $\mathbf{S}_{(r,1^{2})}\oplus\mathbf{S}_{(r-1,2,1)}\oplus\mathbf{S}_{(r-1,1^{3})}.$
In particular, since $r>3$ by hypothesis, $\Lambda^{r+1}\otimes\Lambda^{1}$ and $\mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}$ have no composition factors in common. The analogous statement holds after applying the isomorphism $(\mathrm{Id}\otimes\tau)$ to one of the factors. This implies that the right hand side of (29) is equal to
$\Big{(}(\Lambda^{r+1}\otimes\Lambda^{1})\cap(\mathrm{Id}\otimes\tau)\big{(}\Lambda^{r+1}\otimes\Lambda^{1}\big{)}\Big{)}\ \oplus\ \Big{(}(\mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1})\cap(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(r-1,1^{2})}\otimes\Lambda^{1}\big{)}\Big{)}.$
By Lemmas 6.1 and 6.2 for the respective terms, the latter identifies with
$\Lambda^{r+2}\oplus\mathbf{S}_{(r,1^{2})},$
giving an upper bound for $\ker\overline{\delta}_{r}$. Since this coincides with the lower bound (neglecting the $\mathfrak{S}_{r}$-action) by Proposition 4.20, this completes the proof of the inductive step and hence of the Theorem. ∎
Remark 6.3.
This argument does not go through for $r=3$, since in that case $\Lambda^{4}\otimes\Lambda^{1}$ and $\mathbf{S}_{(2,1^{2})}\otimes\Lambda^{1}$ have a common factor of $\mathbf{S}_{(2,1^{3})}$.